A Bibliography of Publications of Riccardo Fazio

Riccardo Fazio University of Messina

Mathematics Sant’Agata, Messina 98166

Italy

Tel: +39 (0) 90 676 3064 FAX: +39 (0) 90 393502

E-mail: rfazio@na-net.ornl.gov (Internet)

13 October 2017 Version 1.23

Abstract

This bibliography records publications of Riccardo Fazio.

Title word cross-reference

1 [FR10]. 3 [JFA03]. A [Faz01b]. L [Faz01b].

- [Faz01b]. -stability [Faz01b].

13th [BK92]. 14th [Ame94]. 1987 [DG88]. 1990 [EG91]. 1991 [AvdH91, BK92]. 1992 [ITV93, OD93].

2000 [FW01]. 2001 [BBCM03].

4th [BBCM03].

’94 [Ame94].

accuracy [JF04, JF06]. Acireale [ITV93]. action [FI09a, FI14]. Adaptive [Faz09a, JF04, JF06]. advanced [BBCM03, ITV93]. Advection [FI09b, FJ09, FI08a, FJ08]. al [Faz90b]. Algorithm [Faz13a, Faz13b]. algorithms [BK92]. Analysis [FIJ^{+}10, Faz89, FE90, Faz91d, Faz92e, Faz96a, Faz02a, ITV93]. analytical [FI14, ITV93]. Applications [Faz96a, BBCM03, EG91, Faz99a, FW01, Faz99b]. Applied [Ame94, AvdH91, BK92, Faz90c, Faz90e, FI08a, Faz08a, OD93]. Approach [Faz96b, Faz98, Faz94d, Faz95, Faz97b, Faz03b]. approccio [Faz95]. April [OD93]. aquarium [FJ06b]. arbitrary [Faz88a]. area [Faz88a]. aspects [OD93]. Atlanta [Ame94]. axisymmetric [Faz87].

bar [Faz92c]. based [FI10a, FI10b]. bio [FJ06b]. bio-chemical [FJ06b]. Blasius [Faz92a, Faz08b, Faz09b, Faz13a, Faz13b]. Boundary [Faz96b, Faz98, FR10, Faz13c, Faz90a, Faz90d, Faz90c, Faz90e, FE90, Faz91a, Faz91c, Faz91d, Faz92a, Faz92b, Faz92c, Faz92e, FE93a, FE93b, Faz94c, Faz97b, Faz97a, Faz01a, Faz02b, Faz02a, Faz03b, FI10c]. boundary-value [Faz90c]. box [Faz03b]. breaking [Faz88a]. BVPs [Faz02b, Faz03b, FJ14, FJ12].

Capillary [FIJ^{+}10, FIJ^{+}12, CPI^{+}06, CPJ^{+}09, CPI^{+}09, FIJ^{+}07, FI09a, FJ10, FI14]. Catania [ITV93]. Central [FR10]. channel [Faz88a]. chemical [FJ06b]. Class [Faz13c, Faz88b, Faz89]. classes [STFG11]. CLAWPACK [FL03]. Comparison [Faz03a]. complexity [JF04, JF06]. Computation [Ame94]. computational [ITV93, OD93, AvdH91, BK92]. concept [Faz99a, Faz99b]. concepts [Faz96a]. Conditions [FR10]. Conference [BBCM03, EG91, OD93, FW01]. Congress [AvdH91, Ame94, BK92]. conservative [Faz03a]. construction [FI06]. constructive [Faz94d]. continuous [DG88]. continuum [Faz87]. contorno [Faz90b]. Coordinates [FR10]. cross [Faz88a]. cross-sectional [Faz88a].

D [FR10, JFA03]. Deck [FI06]. defined [Faz95]. definiti [Faz95]. deformations [Faz87]. Detection [CCC^{+}10b, CCC^{+}10a]. difference [Faz88b, Faz89, FE93a, FJ12, FJ14]. differential [AvdH91, Faz95]. differenziali [Faz95]. Dimensional [FIJ^{+}12, CPJ^{+}09, CPI^{+}09, FL98, Faz01a, FL03, FI09a, FJ10, FI14]. discontinuity [Faz90f]. Dispersive [FI09b, FJ06a]. Dissipative [FJ91]. domains [Faz95, FI10c]. domini [Faz95]. driven [FIJ^{+}07]. driving [VFM^{+}14]. Dublin [AvdH91, BK92]. Dynamics [FIJ^{+}10, FIJ^{+}12, CPI^{+}06, CPJ^{+}09, CPI^{+}09, FI09a, FJ10, FI14, JFA03].

Effects [FJ91]. eighth [FW01]. elastic [Faz87, Faz88a]. entanglement [STFG11]. Entrapped [FI09a, FI14]. ENUMATH [BBCM03]. Environment [CCC^{+}10b, CCC^{+}10a]. Equation [FJ09, Faz94a, Faz94b, FJ08]. equations [AvdH91, Faz88b, Faz89]. equivalence [Faz93a, Faz93b, FI08b, FI10a, FI10b]. error [FI08a]. estimation [Faz91b, Faz92d]. Eulerian [FR01]. European [BBCM03]. evaluations [Faz90f]. existence [Faz97a]. explicit [Faz10]. Extended [FIJ^{+}10, FIJ^{+}12]. Extension [Faz13a, Faz13b].

failure [Faz10]. Falkner

[Faz94a, Faz94b, Faz13a, Faz13b]. February [FW01]. February/March [FW01]. filled [Faz88a]. Finite [Faz89, FJ12, FJ14, Faz87, Faz88b]. first [FI08a]. flow [Faz91b, Faz92d]. flows [FIJ^{+}07]. fluid [Faz88a]. Flux [FJ09]. formulated [Faz92a]. formulation [FI10c]. Fourth [OD93]. Free [Faz98, Faz90a, Faz90d, Faz90e, FE90, Faz91a, Faz92a, Faz92b, Faz97b, Faz97a, Faz02b, Faz03b].

gas [FI09a, FI14]. Gathering [FI06]. Georgia [Ame94]. grids [FJ12, FJ14]. Group [Faz94a, Faz89, Faz94b, ITV93]. Groups [FI10d, FI10a, FI10b].

Horizontal [FIJ^{+}10, FIJ^{+}12, CPI^{+}06, CPJ^{+}09, CPI^{+}09, FJ10]. Hyperbolic [EG91, Faz88b, Faz89, Faz90f, Faz90d, Faz91d, Faz92c, FE93a, FE93b, FL98, FR01, Faz02a, Faz03a, FL03, OD93, FW01].

identification [Faz02b]. II [AvdH91]. Ill [FJ10]. IMACS [AvdH91, Ame94, BK92, Ame94]. immiscible [FIJ^{+}07]. impact [Faz92c]. Implicit [FJ06a, FE93a, FI08a]. Industrial [CCC^{+}10b, CCC^{+}09, CCC^{+}10a]. Infinite [Faz96b, Faz95, Faz97b, Faz02b, Faz03b, FJ12, FJ14]. infiniti [Faz95]. initial [Faz01b]. initial-value [Faz01b]. instance [Faz10]. Institute [Ame94]. integrazione [Faz90b]. Interface [FR10, FR01, Faz03a]. International [EG91, OD93, FW01, ITV93]. Intervals [Faz96b, Faz97b, Faz02b, Faz03b, FJ12, FJ14]. Invariance [Faz87, FIJ^{+}12, Faz13c, Faz94a, Faz94b, Faz08a]. Ireland [AvdH91, BK92]. Ischia [BBCM03]. isotropic [STFG11]. Italian [Faz95]. Italy [ITV93, OD93]. Iterative [Faz12, Faz13c, Faz91a, Faz91b, Faz92b, Faz93a, Faz94c, Faz01a, FI10a, FI10b, FI10d, Faz15]. IVPs [Faz09a].

July [AvdH91, Ame94, BK92, BBCM03]. June [EG91].

Keller [Faz03b].

Lagrangian [FR01, FR10]. length [Faz91b, Faz92d]. libero [Faz90b]. Limiters [FJ09]. Linear [FI09b, Faz90f, FI08a]. Liquid [CCC^{+}09, FIJ^{+}10, FIJ^{+}12, CPJ^{+}09, CPI^{+}09, FJ10, FI14]. liquids [CPI^{+}06, FIJ^{+}07]. Local [FI08a, STFG11]. low [JF04, JF06].

Magdeburg [FW01]. Magic [FI06]. material [Faz92c]. materials [Faz91c]. Mathematical [CPI^{+}06, FJ06b, CPJ^{+}09, CPI^{+}09, ITV93, JFA03]. Mathematics [Ame94, AvdH91, BK92, BBCM03]. MATLAB [Faz10]. means [FJ09]. measurements [STFG11]. media [DG88]. mesh [FL98, FL03]. Method [Faz12, Faz13c, Faz90a, Faz90c, Faz90e, Faz91a, Faz91b, Faz92b, Faz94c, Faz01a, Faz15]. Methods [FI09b, FI10d, EG91, Faz88b, Faz89, Faz93a, Faz93b, Faz94d, FL98, Faz01b, FL03, FJ06a, FI08a, Faz08b, Faz09b, FI10a, FI10b, ITV93]. Metodi [Faz90b]. mobile [Faz90b]. model [FIJ^{+}07, Faz08a, Faz13a, Faz13b, JFA03]. modeling [CPI^{+}06, CPJ^{+}09, CPI^{+}09, FJ06b]. Models [FIJ^{+}12, FI08a, FI09a, FJ10]. Modern [ITV93]. Moving [FL98, FL03, Faz13c, Faz91c, Faz91d, Faz92b, Faz92c, Faz92e, FE93a, FE93b, Faz94c, Faz01a, Faz02a, FI10c]. Moving-mesh [FL98, FL03]. mucilage [JFA03].

Non [Faz93a, FI10d, Faz91a, FI10a, FI10b]. Non-iterative [Faz93a, FI10d, Faz91a, FI10a, FI10b]. Noniterative [Faz93b, Faz90a, Faz90c]. Nonlinear [FJ91, Faz90d, OD93]. Normal [Faz90e]. Novel [Faz96b]. numerica [Faz90b]. Numerical [BBCM03, Faz90f, Faz92d, FE93b, Faz94d, Faz96b, Faz98, Faz99a, Faz08a, Faz09b, Faz99b, CPI^{+}06, CPJ^{+}09, CPI^{+}09, EG91, FE90, Faz94a, Faz94b, Faz94c, Faz95, Faz96a, Faz97a, Faz01a, Faz01b, Faz02b, Faz02a, FJ06b, FI08b, Faz09a, FI14, BBCM03]. numerico [Faz95]. numerics [FW01].

o [Faz90b]. October [DG88, ITV93]. ODE45 [Faz10]. One [CPJ^{+}09, CPI^{+}09, FIJ^{+}12, FL98, Faz01a, FL03, FI09a, FJ10, FI14]. One-Dimensional [FIJ^{+}12, CPJ^{+}09, CPI^{+}09, FL98, Faz01a, FL03, FI09a, FJ10, FI14]. open [VFM^{+}14]. Order [FJ09, FR10, FI08a].

papers [AvdH91, BK92]. Parabolic [Faz13c, Faz92e, Faz94c, Faz01a, FI10c]. March [FW01]. Penetrant [CCC^{+}10b, CCC^{+}09, CCC^{+}10a]. penetration [FIJ^{+}07]. physics [ITV93]. Piston [FR10]. point [Faz90c]. Pol [Faz08a]. posed [FJ10]. Positive [FJ08, FJ09]. potentials [Faz87]. prediction [JFA03]. Probability [CCC^{+}10a, CCC^{+}10b]. Problem [Faz12, Faz13a, Faz13b, Faz90a, Faz90d, Faz91a, Faz91c, Faz91d, Faz92a, Faz92c, FE93a, FE93b, Faz97b, Faz02a, Faz08b, Faz09b, Faz15]. problemi [Faz90b, Faz95]. Problems [EG91, Faz96b, Faz98, FI09b, FR10, Faz13c, Faz90c, Faz90e, FE90, Faz92b, Faz92e, Faz94c, Faz95, Faz97b, Faz97a, FL98, Faz01a, FR01, Faz01b, Faz03a, FL03, FJ06a, FI10c, FW01, OD93]. proceedings [Ame94, BBCM03, EG91, ITV93, OD93]. process [CCC^{+}09]. propagation [FJ06a]. properties [Faz87]. Pseudo [FI09b, FJ06a, FI08a]. Pseudo-Spectral [FI09b, FJ06a, FI08a].

Quantitative [STFG11]. quantum [VFM^{+}14]. quasi [Faz90f, FJ12, FJ14]. quasi-linear [Faz90f]. quasi-uniform [FJ12, FJ14]. quasilinear [Faz88b, Faz89]. quasilinear-hyperbolic [Faz88b, Faz89].

rate [Faz91c, Faz92c]. rate-type [Faz91c, Faz92c]. reactors [Faz91b, Faz92d]. reduction [FI08a]. related [Faz90f]. remarks [Faz88b]. revised [AvdH91, BK92].

Sakiadis [Faz12, Faz15]. Scaling [FI08b, FIJ^{+}12, Faz13c, Faz99a, Faz08a, FI10a, FI10b, Faz99b]. scheme [FE93a, Faz03b]. Schemes [FJ09, FR10, FR01, Faz03a, FJ08, FJ12, FJ14]. Second [FJ09, FR10]. sectional [Faz88a]. selected [AvdH91, BK92]. shock [Faz90f]. shooting [Faz90a, Faz91a]. Similarity [FE90, Faz91d, Faz92e, Faz98, Faz02a, FIJ^{+}10, Faz90f, Faz96a]. singular [Faz02a]. Skan [Faz94a, Faz94b, Faz13a, Faz13b]. Solution [Faz96b, Faz98, FE93b, Faz94a, Faz94c, Faz97a, Faz01a]. solutions [Faz90f, Faz94b]. solver [Faz10]. solvers [Faz09a, JF04, JF06]. Some [Faz88b]. Spectral [FI09b, FJ06a, FI08a]. spiral [FI08b, FI10a, FI10b]. stability [DG88, Faz01b]. stiff [JF04, JF06].

Stiffness [Faz01b]. strategies

[FI06, Faz09a]. stress [Faz92c]. study [FI14]. surface [FIJ^{+}07]. surface-driven [FIJ^{+}07]. survey [Faz02b]. Sweden [EG91]. systems [Faz90f, VFM^{+}14].

Taormina [DG88, OD93]. Technology [Ame94]. test [Faz90a, Faz91a, Faz97b, Faz97a]. Testing [CCC^{+}10b, CCC^{+}09, CCC^{+}10a]. theoretical [OD93]. theory [BK92, EG91, Faz94b, FW01]. Third [EG91]. Töpfer [Faz13b, Faz13a]. Transformation [Faz08b, FI10d, Faz12, Faz13c, Faz90a, Faz90c, Faz90e, Faz91a, Faz91b, Faz92b, Faz93a, Faz93b, Faz94c, Faz94d, Faz01a, Faz09b, FI10a, FI10b, Faz15]. Transitionless [VFM^{+}14]. Translation [FI10d]. trasformazione [Faz90b]. truncated [Faz02b]. tube [Faz88a]. tubular [Faz91b, Faz92d]. two [Faz90c, Faz03a, FIJ^{+}07]. two-point [Faz90c]. type [Faz91c, Faz92c].

unbounded [FI10c]. uniform [FJ12, FJ14]. uniqueness [Faz97a]. Uppsala [EG91]. USA [Ame94]. using [FL03].

Value [Faz96b, Faz90d, Faz90c, Faz90e, FE90, Faz92a, Faz92b, Faz97b, Faz01b]. variables [Faz90e]. variants [Faz08b, Faz09b]. velocity [Faz90f]. via [STFG11]. viewpoint [FI08b].

wave [FJ06a]. Waves [FJ91, Faz88a, DG88]. well [FJ10]. well-posed [FJ10]. Werner [STFG11]. within [Faz94b]. witnesses [STFG11]. workshop [ITV93]. World [AvdH91, Ame94, BK92].

References

Ames:1994:IPI

[Ame94] William F. Ames. IMACS ’94:

proceedings of the 14th IMACS World Congress on Computation and Applied Mathematics: July 11–15, 1994, Georgia Institute of Technology, Atlanta, Georgia, USA. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1994. 1572 pp. Three volumes.

Ames:1991:CAM

[AvdH91] W. F. Ames and P. J. van der Houwen, editors. Computational and applied mathematics, II: differential equations: selected and revised papers from the IMACS World Congress, Dublin, Ireland, July 1991. North-Holland, Amsterdam, The Netherlands, 1991. ISBN 0-444-89702-X. LCCN QA370 .C625 1991.

Brezzi:2003:NMA

[BBCM03] Franco Brezzi, Annalisa Buffa, Stefania Corsaro, and Almerico Murli, editors. Numerical mathematics and advanced applications: proceedings of ENUMATH 2001, the 4th European Conference on Numerical Mathematics and Advanced Applications, Ischia, July 2001. Springer, Milano, Italy, 2003. ISBN 88-470-0180-3. LCCN QA297 .E89 2001.

Brezinski:1992:CAM

[BK92] C. Brezinski and U. Kulisch, editors. Computational and applied mathematics, I: algorithms and theory: selected and revised papers from the IMACS 13th World Congress, Dublin, Ireland, July 1991. North-Holland, Amsterdam, The Netherlands, 1992. LCCN QA76.95 .C62 1991.

Caturano:2009:LPT

[CCC^{+}09] Gennaro Caturano, Giovanni Cavaccini, Antonio Ciliberto, Vittoria Pianese, and Riccardo Fazio. Liquid penetrant testing: industrial process. Comm. SIMAI Congress, 3:319.1–319.12, 2009.

Caturano:2010:PDP

[CCC^{+}10a] Gennaro Caturano, Giovanni Cavaccini, Antonio Ciliberto, Vittoria Pianese, and Riccardo Fazio. Probability of detection for penetrant testing in industrial environment. In Applied and industrial mathematics in Italy III, volume 82 of Ser. Adv. Math. Appl. Sci., pages 186–195. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010.

Caturano:2010:POD

[CCC^{+}10b] Gennaro Caturano, Giovanni Cavaccini, Antonio Ciliberto, Vittoria Pianese, and Riccardo Fazio. Probability of detection for penetrant testing in industrial environment. In E. De Bernardis, R. Spligher, and V. Valenti, editors, Applied and Industrial Mathematics in Italy III, pages 186–195. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010.

Cavaccini:2006:MNM

[CPI^{+}06] Giovanni Cavaccini, Vittoria Pianese, Salvatore Iacono, Alessandra Jannelli, and Riccardo Fazio. Mathematical and numerical modeling of liquids dynamics in a horizontal capillary. In T. Simos and G. Maroulis, editors, Recent Progress in Computational Sciences and Engineering, ICCMSE 2006, Lecture Series on Computer and Computational Sciences, volume 7, pages 66–70. Koninklijke Brill NV, Leiden, The Netherlands, 2006.

Cavaccini:2009:OMN

[CPI^{+}09] Giovanni Cavaccini, Vittoria Pianese, Salvatore Iacono, Alessandra Jannelli, and Riccardo Fazio. One-dimensional mathematical and numerical modeling of liquid dynamics in a horizontal capillary. J. Comput. Meth. Sci. Eng., 9:3–16, 2009.

Cavaccini:2009:ODM

[CPJ^{+}09] Giovanni Cavaccini, Vittoria Pianese, Alessandra Jannelli, Salvatore Iacono, and Riccardo Fazio. One-dimensional mathematical and numerical modeling of liquid dynamics in a horizontal capillary. Journal of Computational Methods in Sciences and Engineering, 9(1-2):3–16, 2009. CO-DEN JCMSCR. ISSN 1472-7978.

Donato:1988:WSC

[DG88] A. Donato and S. Giambó, editors. Waves and stability in continuous media, Taormina, October 1987. Editel, Commenda di Rende, Italy, 1988. LCCN QA927 .W39 1988. Papers presented to the IVth biannual Meeting on Waves and Stability in Continuous Media held in Taormina (Italy), October 5–10, 1987.

Engquist:1991:TIC

[EG91] Bjorn Engquist and Bertil Gustafsson. Third International Conference on Hyperbolic Problems: theory, numerical methods and applications: proceedings, Uppsala, Sweden, June 11–15, 1990. Chartwell-Bratt, Bromley, Kent, England, 1991. ISBN 91-44-33421-4 (Studentlitteratur), 0-86238-285-8 (Chartwell Bratt). LCCN QA377.I563 1990. The two earlier conferences … were held in St. Etienne, France, 1986 and Aachen, Germany, 1988.

Fazio:1987:IPE

[Faz87] Riccardo Fazio. Invariance properties and elastic potentials for an axisymmetric continuum with finite deformations. Atti Sem. Mat. Fis. Univ. Modena, XXXV:147– 158, 1987.

Fazio:1988:BWC

[Faz88a] Riccardo Fazio. On the breaking waves in a channel or in a fluid filled elastic tube with arbitrary cross-sectional area. Riv. Mat. Univ. Parma, 14:71–79, 1988. ISSN 0035-6298.

Fazio:1988:SRA

[Faz88b] Riccardo Fazio. Some remarks about finite difference methods for a class of quasilinearhyperbolic equations. In Donato and Giambó [DG88], pages 115–119. LCCN QA927 .W39 1988. Papers presented to the IVth biannual Meeting on Waves and Stability in Continuous Media held in Taormina (Italy), October 5–10, 1987.

Fazio:1989:FDM

[Faz89] Riccardo Fazio. Finite difference methods and group analysis for a class of quasilinear-hyperbolic equations. Int. J. Computer Math., 28:151–160, 1989. CO-DEN IJCMAT. ISSN 0020-7160.

Fazio:1990:FBT

[Faz90a] Riccardo Fazio. A free boundary test problem for a noniterative transformation method and a shooting method. Atti Accad. Peloritana Pericolanti Cl. Sci. Fis. Mat. Natur., 68(suppl. 1):141–151 (1991), 1990. CO-DEN AAPFAO. ISSN 0365-0359. Memoirs of mathematical physics in honor of Giovanni Carini on his 70th birthday (Italian).

Fazio:1990:MTP

[Faz90b] Riccardo Fazio. Metodi di trasformazione per l’integrazione numerica di problemi al contorno libero o mobile. Ph.d. thesis, Universitá di Messina, Messina, Italy, 1990.

Fazio:1990:NTM

[Faz90c] Riccardo Fazio. A noniterative transformation method applied to two-point boundary-value problems. Appl. Math. Comput., 39 (1):79–87, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fazio:1990:NHF

[Faz90d] Riccardo Fazio. A nonlinear hyperbolic free boundary value problem. Acta Mech., 81(3–4): 221–226, 1990. CODEN AMH-CAP. ISSN 0001-5970.

Fazio:1990:NVT

[Faz90e] Riccardo Fazio. Normal variables transformation method applied to free boundary value problems. Int. J. Computer Math., 37:189–199, 1990. CODEN IJC-MAT. ISSN 0020-7160.

Fazio:1990:NES

[Faz90f] Riccardo Fazio. Numerical evaluations for shock and discontinuity velocity related to similarity solutions of quasi-linear hyperbolic systems. Journal of Computational and Applied Mathematics, 30(3):341–349, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fazio:1991:FBT

[Faz91a] Riccardo Fazio. A free boundary test problem for a noniterative transformation method and a shooting method. Atti Accad. Peloritana Pericolanti Cl. Sci. Fis. Mat. Natur., LXVIII: 141–151, 1991.

Fazio:1991:ITM

[Faz91b] Riccardo Fazio. The iterative transformation method and length estimation for tubular flow reactors. Appl. Math. Comput., 42(2 (part II)):105–110, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fazio:1991:MBP

[Faz91c] Riccardo Fazio. A moving boundary problem for rate-type materials. In Fusco and Jeffrey [FJ91], pages 84–90. ISBN 0-582-08732-5. ISSN 0269-3674. LCCN QA927 .N667 1991; Math.

Fazio:1991:SAM

[Faz91d] Riccardo Fazio. Similarity analysis of a moving boundary hyperbolic problem. In Third International Conference on Hyperbolic Problems, Vol. I, II (Uppsala, 1990) [EG91], pages 409–417. ISBN 91-44-33421-4 (Studentlitteratur), 0-86238-285-8 (Chartwell Bratt). LCCN QA377.I563 1990. The two earlier conferences … were held in St. Etienne, France, 1986 and Aachen, Germany, 1988.

Fazio:1992:BPF

[Faz92a] Riccardo Fazio. The Blasius problem formulated as a free boundary value problem. Acta Mech., 95(1-4):1–7, 1992. CODEN AMHCAP. ISSN 0001-5970.

Fazio:1992:ITM

[Faz92b] Riccardo Fazio. The iterative transformation method: free or moving boundary value problems. In Brezinski and Kulisch [BK92], pages 147–152. LCCN QA76.95 .C62 1991.

Fazio:1992:MBH

[Faz92c] Riccardo Fazio. A moving boundary hyperbolic problem for a stress impact in a bar of ratetype material. Wave Motion, 16(4):299–305, 1992. CODEN WAMOD9. ISSN 0165-2125.

Fazio:1992:NLE

[Faz92d] Riccardo Fazio. Numerical length estimation for tubular flow reactors. Journal of Computational and Applied Mathematics, 41(3):313–321, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fazio:1992:SAM

[Faz92e] Riccardo Fazio. Similarity analysis for moving boundary parabolic problems. In Ames and van der Houwen [AvdH91], pages 153– 162. ISBN 0-444-89702-X. LCCN QA370 .C625 1991.

Fazio:1993:NIT

[Faz93a] Riccardo Fazio. Non-iterative transformation methods equivalence. In Ibragimov et al. [ITV93], pages 217–221. ISBN 0-7923-2480-3. LCCN QC19.2 .M63 1992. Papers presented at the workshop on Modern Group Analysis, held at Catania in Sicily, Italy, 1992.

Fazio:1993:NTM

[Faz93b] Riccardo Fazio. Noniterative transformation methods equivalence. In Modern group analysis: advanced analytical and computational methods in mathematical physics (Acireale, 1992), pages 217–221. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993.

Fazio:1994:FSEa

[Faz94a] Riccardo Fazio. The Falkner– Skan equation: Group invariance and numerical solution, 1994. Three volumes.

Fazio:1994:FSEb

[Faz94b] Riccardo Fazio. The Falkner– Skan equation: numerical solutions within group invariance theory. Calcolo, 31(1–2):115– 124, 1994. CODEN CALOBK. ISSN 0008-0624 (print), 1126-5434 (electronic).

Fazio:1994:ITM

[Faz94c] Riccardo Fazio. The iterative transformation method: numerical solution of moving boundary parabolic problems, 1994. Three volumes.

Fazio:1994:NTM

[Faz94d] Riccardo Fazio. Numerical transformation methods: a constructive approach. Journal of Computational and Applied Mathematics, 50(1–3):299–303, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fazio:1995:ANP

[Faz95] Riccardo Fazio. Un approccio numerico per problemi differenziali definiti su domini infiniti. (Italian) [a numerical approach to differential problems defined on infinite domains], 1995. Atti del XV Congresso dell’Unione Matematica Italiana, Padova, 11–16 Settembre 1995.

Fazio:1996:ASC

[Faz96a] Riccardo Fazio. Applications of similarity concepts to numerical analysis, 1996. Final program of the 1996 SIAM Annual Meeting, Kansas City, July 22–26 1996.

Fazio:1996:NAN

[Faz96b] Riccardo Fazio. A novel approach to the numerical solution of boundary value problems on infinite intervals. SIAM J. Numer. Anal., 33(4):1473–1483, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        toc/SINUM/33/4.

Fazio:1997:NTE

[Faz97a] Riccardo Fazio. A numerical test for the existence and uniqueness of solution of free boundary problems. Applicable Analysis, 66(1-2):89–100, 1997. CO-DEN APANCC. ISSN 0003-6811.

Fazio:1997:ATP

[Faz97b] Riccardo Fazio. A test problem for a free boundary approach to boundary value problems on infinite intervals, 1997. Final program of the SciCADE97 International Conference, Grado, September 15–19 1997.

Fazio:1998:SAN

[Faz98] Riccardo Fazio. A similarity approach to the numerical solution of free boundary problems. SIAM Rev., 40(3):616– 635, September 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fazio:1999:NAS

[Faz99a] Riccardo Fazio. Numerical applications of the scaling concept. Acta Applicandae Mathematicae, 55:1–25, 1999. CODEN AA-MADV. ISSN 0167-8019.

MR1675478

[Faz99b] Riccardo Fazio. Numerical applications of the scaling concept. Acta Applicandae Mathematicae, 55(1):1–25, 1999. CODEN AA-MADV. ISSN 0167-8019.

Fazio:2001:ITM

[Faz01a] Riccardo Fazio. The iterative transformation method: numerical solution of one-dimensional parabolic moving boundary problems. Int. J. Computer Math., 78 (2):213–223, 2001. CODEN IJC-MAT. ISSN 0020-7160.

Fazio:2001:SNI

[Faz01b] Riccardo Fazio. Stiffness in numerical initial-value problems: A- and L-stability of numerical methods. International journal of mathematical education in science and technology, 32(5):752– 760, 2001. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Fazio:2002:SNA

[Faz02a] Riccardo Fazio. Similarity and numerical analysis of a singular moving boundary hyperbolic problem. In A. M. Anile, V. Capasso, and A. Greco, editors, Progress in industrial mathematics at ECMI 2000 (Palermo), volume 1 of Math. Ind., pages 339– 344. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002.

Fazio:2002:SFB

[Faz02b] Riccardo Fazio. A survey on free boundary identification of the truncated boundary in numerical BVPs on infinite intervals. Journal of Computational and Applied Mathematics, 140(1–2):331– 344, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fazio:2003:CTC

[Faz03a] Riccardo Fazio. Comparison of two conservative schemes for hyperbolic interface problems. In Brezzi et al. [BBCM03], pages 85– 93. ISBN 88-470-0180-3. LCCN QA297 .E89 2001.

Fazio:2003:FBA

[Faz03b] Riccardo Fazio. A free boundary approach and Keller’s box scheme for BVPs on infinite intervals. Int. J. Computer Math., 80(12):1549– 1560, 2003. CODEN IJCMAT. ISSN 0020-7160.

Fazio:2008:NSI

[Faz08a] Riccardo Fazio. Numerical scaling invariance applied to the van der Pol model. Acta Applicandae Mathematicae, 104(1):107– 114, 2008. CODEN AAMADV. ISSN 0167-8019.

Fazio:2008:TMB

[Faz08b] Riccardo Fazio. Transformation methods for the Blasius problem and its recent variants. In S. I. Ao, L. Gelman, D. W. L. Hukins, A. Hunter, and A. M. Korsunsky, editors, Proceedings World Congress on Engineering 2008, held in London, 2–4 July, volume II, pages 938–943. IAENG, Kwun Tong — Hong Kong, 2008.

Fazio:2009:ASN

[Faz09a] Riccardo Fazio. Adaptive strategies for numerical IVPs solvers. In S. I. Ao, L. Gelman, D. W. L. Hukins, A. Hunter, and A. M. Korsunsky, editors, Proceedings World Congress on Engineering 2009, held in London, 1–3 July, volume II, pages 1213–1218. IAENG, Kwun Tong — Hong Kong, 2009.

Fazio:2009:NTM

[Faz09b] Riccardo Fazio. Numerical transformation methods: Blasius problem and its variants. Appl. Math. Comput., 215(4):1513– 1521, October 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fazio:2010:IFM

[Faz10] Riccardo Fazio. An instance of failure for the MATLAB explicit ODE45 solver. In S. I. Ao, L. Gelman, D. W. L. Hukins, A. Hunter, and A. M. Korsunsky, editors, Proceedings World Congress on Engineering 2010, held in London, 30 June–2 July, volume III, pages 1935–1940. IAENG, Kwun Tong — Hong Kong, 2010.

Fazio:2012:ITM

[Faz12] Riccardo Fazio. The iterative transformation method for the Sakiadis problem, 2012. Preprint available at the URL: mat521.

                                        unime.it/fazio/preprints/Sakiadis.
                                        pdf.

Fazio:2013:BPFa

[Faz13a] Riccardo Fazio. Blasius problem and Falkner–Skan model: Töpfer’s algorithm and its exten-

                                        sion, 2013. URL http://arxiv.
                                        org/pdf/1212.5057v1.pdf.

Fazio:2013:BPFb

[Faz13b] Riccardo Fazio. Blasius problem and Falkner–Skan model: Töpfer’s algorithm and its extension. Computers and Fluids, 73:202–209, 2013. URL http:/

                                        /arxiv.org/pdf/1212.5057v1.
                                        pdf.

Fazio:2013:SII

[Faz13c] Riccardo Fazio. Scaling invariance and the iterative transformation method for a class of parabolic moving boundary problems. Int. J. Non-Linear Mech., 50:136–140, 2013. Preprint avail-

                                        able at http://arxiv.org/pdf/
                                        1211.6703v1.pdf.

Fazio:2015:ITM

[Faz15] Riccardo Fazio. The iterative transformation method for the Sakiadis problem. Computers and Fluids, 106:196–200, 2015. CO-DEN CPFLBI. ISSN 0045-7930 (print), 1879-0747 (electronic).

Fazio:1990:SNA

[FE90] Riccardo Fazio and David J.

Evans. Similarity and numerical analysis for free boundary value problems. Int. J. Computer Math., 31:215–220, 1990. CO-DEN IJCMAT. ISSN 0020-7160. 39:249, 1991.

Fazio:1993:IDS

[FE93a] Riccardo Fazio and David J.

Evans. An implicit difference scheme for a moving boundary hyperbolic problem. Applied Numerical Mathematics: Transactions of IMACS, 12(6):485–496, August 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fazio:1993:NSM

[FE93b] Riccardo Fazio and David J.

Evans. Numerical solution of a moving boundary hyperbolic problem. In Oliveri and Donato [OD93], pages 217–226. ISBN 3-528-07643-7. ISSN 0179-9614. LCCN QA377.I565 1992.

Fazio:2006:DCS

[FI06] Riccardo Fazio and Salvatore Iacono. Deck construction strategies for Magic: the Gathering, 2006. VIII Congresso SIMAI, Baia Samuele, 22-26 Maggio 2006, 6 pp.

Fazio:2008:LER

[FI08a] Riccardo Fazio and Salvatore Iacono. Local error reduction for first order implicit pseudospectral methods applied to linear advection models. In N. Manganaro, R. Monaco, and S. Rionero, editors, Proceedings “WASCOM 2007” 14th Conference on Waves and Stability in Continuous Media, pages 258–263. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008.

Fazio:2008:SSE

[FI08b] Riccardo Fazio and Salvatore Iacono. Scaling and spiral equivalence from a numerical viewpoint. In S. I. Ao, L. Gelman, D. W. L. Hukins, A. Hunter, and A. M. Korsunsky, editors, Proceedings World Congress on Engineering 2008, held in London, 2–4 July, volume II, pages 1015–1020. IAENG, Kwun Tong — Hong Kong, 2008.

Fazio:2009:EGA

[FI09a] Riccardo Fazio and Salvatore Iacono. Entrapped gas action for one-dimensional models of capillary dynamics. In S. I. Ao, L. Gelman, D. W. L. Hukins, A. Hunter, and A. M. Korsunsky, editors, Proceedings World Congress on Engineering 2009, held in London, 1–3 July, volume II, pages 1016–1021. IAENG, Kwun Tong — Hong Kong, 2009.

Fazio:2009:PSM

[FI09b] Riccardo Fazio and Salvatore Iacono. Pseudo-spectral methods for linear advection and dispersive problems. IAENG Inter. J.

Appl. Math., 39(1):36–47, 2009. ISSN 1992-9978.

Fazio:2010:ENI

[FI10a] Riccardo Fazio and Salvatore Iacono. On the equivalence of noniterative transformation methods based on scaling and spiral groups. Math. Methods Appl. Sci., 33(5):585–591, 2010. ISSN 0170-4214.

Fazio:2010:ESI

[FI10b] Riccardo Fazio and Salvatore Iacono. On the equivalence of noniterative transformation methods based on scaling and spiral groups. Math. Meth. Appl. Sci., 33:585–591, 2010.

Fazio:2010:MBF

[FI10c] Riccardo Fazio and Salvatore Iacono. On the moving boundary formulation for parabolic problems on unbounded domains. Int. J. Computer Math., 87(1-3):186– 198, 2010. CODEN IJCMAT. ISSN 0020-7160.

Fazio:2010:TGN

[FI10d] Riccardo Fazio and Salvatore Iacono. On the translation groups and non-iterative transformation methods. In E. De Bernardis, R. Spligher, and V. Valenti, editors, Applied and Industrial Mathematics in Italy III, Ser. Adv. Math. Appl. Sci., pages 331– 340. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010.

Fazio:2014:ANS

[FI14] Riccardo Fazio and Salvatore Iacono. An analytical and numerical study of liquid dynamics in a one-dimensional capillary under entrapped gas action. Mathematical Methods in the Applied Sciences, 37(18):2923–2933, 2014. CODEN MMSCDB. ISSN 0170-4214 (print), 1099-1476 (electronic).

Fazio:2007:TIL

[FIJ^{+}07] Riccardo Fazio, Salvatore Iacono, Alessandra Jannelli, Giovanni Cavaccini, and Vittoria Pianese. A two immiscible liquids penetration model for surfacedriven capillary flows. Proceedings in Applied Mathematics and Mechanics, 7:2150003– 2150004, 2007. ISSN 1617-7061.

Fazio:2010:LDH

[FIJ^{+}10] Riccardo Fazio, Salvatore Iacono, Alessandra Jannelli, Giovanni Cavaccini, and Vittoria Pianese. Liquid dynamics in a horizontal capillary: Extended similarity analysis. In E. De Bernardis, R. Spligher, and V. Valenti, editors, Applied and Industrial Mathematics in Italy III, Ser. Adv. Math. Appl. Sci., pages 341– 352. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010.

Fazio:2012:ESI

[FIJ^{+}12] Riccardo Fazio, Salvatore Iacono, Alessandra Jannelli, Giovanni Cavaccini, and Vittoria Pianese. Extended scaling invariance of one-dimensional models of liquid dynamics in a horizontal capillary. Math. Methods Appl. Sci., 35(8):935–942, 2012. ISSN 0170-4214.

Fusco:1991:NWD

[FJ91] Domenico Fusco and Alan Jeffrey, editors. Nonlinear Waves and Dissipative Effects, volume 227 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, 1991. ISBN 0-582-08732-5. ISSN 0269-3674. LCCN QA927 .N667 1991; Math.

Fazio:2006:IPS

[FJ06a] Riccardo Fazio and Alessandra Jannelli. Implicit pseudo-spectral methods for dispersive and wave propagation problems, 2006. VIII Congresso SIMAI, Baia Samuele, 22-26 Maggio 2006, 6 pp.

Fazio:2006:MNM

[FJ06b] Riccardo Fazio and Alessandra Jannelli. Mathematical and numerical modeling for a bio-chemical aquarium. Appl. Math. Comput., 174(2):1370– 1383, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fazio:2008:PSA

[FJ08] Riccardo Fazio and Alessandra Jannelli. Positive schemes for the advection equation. In Natale Manganaro, Roberto Monaco, and Salvatore Rionero, editors, Proceedings “WASCOM 2007”

14th Conference on Waves and Stability in Continuous Media, pages 264–273. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008.

Fazio:2009:SOP

[FJ09] Riccardo Fazio and Alessandra Jannelli. Second order positive schemes by means of flux limiters for the advection equation. IAENG Inter. J. Appl. Math., 39 (1):25–35, 2009. ISSN 1992-9978.

Fazio:2010:IWP

[FJ10] Riccardo Fazio and Alessandra Jannelli. Ill and well-posed onedimensional models of liquid dynamics in a horizontal capillary. In E. De Bernardis, R. Spligher, and V. Valenti, editors, Applied and industrial mathematics in Italy III, volume 82 of Ser. Adv. Math. Appl. Sci., pages 353–364. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010.

Fazio:2012:FDS

[FJ12] Riccardo Fazio and Alessandra Jannelli. Finite difference schemes on quasi-uniform grids for Bvps on infinite intervals, 2012. Preprint available at http:

                                        //arxiv.org/abs/1211.5427.

Fazio:2014:FDS

[FJ14] Riccardo Fazio and Alessandra Jannelli. Finite difference schemes on quasi-uniform grids for BVPs on infinite intervals. Journal of Computational and Applied Mathematics, 269:14– 23, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fazio:1998:MMM

[FL98] Riccardo Fazio and Randall J.

LeVeque. Moving-mesh methods for one-dimensional hyperbolic problems. In Abstracts of the SIMAI98 Conference, volume 1, pages 314–317. ????, ????, 1998. Giardini Naxos, June 1–5 1998.

Fazio:2003:MMM

[FL03] Riccardo Fazio and Randall J.

LeVeque. Moving-mesh methods for one-dimensional hyperbolic problems using CLAW-PACK. Computers and Mathematics with Applications, 45(1-3): 273–298, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). Numerical methods in physics, chemistry, and engineering.

Fazio:2001:LCS

[FR01] Riccardo Fazio and Giovanni Russo. Eulerian and Lagrangian schemes for hyperbolic interface problems. In Freistühler and Warnecke [FW01], pages 347– 356. ISBN 3-7643-6711-3 (set), 3-7643-6709-1 (vol. 1), 3-7643-6710-5 (vol. 2). LCCN QA377 .I575 2001; QA377 .I565 2000. Two volumes.

Fazio:2010:CSS

[FR10] Riccardo Fazio and Giovanni Russo. Central schemes and second order boundary conditions for 1D interface and piston problems in Lagrangian coordinates. Communications in Computational Physics, 8(4):797–822, 2010. ISSN 1815-2406 (print), 1991-7120 (electronic).

Freistuhler:2001:HPT

[FW01] Heinrich Freistühler and Gerald Warnecke, editors. Hyperbolic problems: theory, numerics, applications: eighth international conference in Magdeburg, February/March 2000, volume 140/ 141 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2001. ISBN 3-7643-6711-3 (set), 3-7643-6709-1 (vol. 1), 3-7643-6710-5 (vol. 2). LCCN QA377 .I575 2001; QA377 .I565 2000. Two volumes.

Ibragimov:1993:MGA

[ITV93] N. Kh. Ibragimov, M. Torrisi, and A. Valenti, editors. Modern group analysis: advanced analytical and computational methods in mathematical physics: proceedings of the international workshop, Acireale, Catania, Italy, October 27–31, 1992. Kluwer, Amsterdam, 1993. ISBN 0-7923-2480-3. xi + 393 pp. LCCN QC19.2 .M63 1992. Papers presented at the workshop on Modern Group Analysis, held at Catania in Sicily, Italy, 1992.

Jannelli:2004:ASS

[JF04] Alessandra Jannelli and Riccardo Fazio. Adaptive stiff solvers at low accuracy and complexity. In T. Simos and G. Maroulis, editors, Proceedings of the ICCMSE 2004, volume 1, pages 230–233. VSP/Brill Academic Publishers, Boston, MA, USA, 2004.

Jannelli:2006:ASS

[JF06] Alessandra Jannelli and Riccardo Fazio. Adaptive stiff solvers at low accuracy and complexity. Journal of Computational and Applied Mathematics, 191(2):246–258, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Jannelli:2003:MMP

[JFA03] Alessandra Jannelli, Riccardo Fazio, and Davide Ambrosi. A 3D mathematical model for the prediction of mucilage dynamics. Computers and Fluids, 32:47–57, 2003. CODEN CPFLBI. ISSN 0045-7930.

Oliveri:1993:NHP

[OD93] Francesco Oliveri and Andrea Donato, editors. Nonlinear hyperbolic problems: theoretical, applied, and computational aspects: proceedings of the Fourth International Conference on Hyperbolic Problems, Taormina, Italy, April 3–8, 1992, volume 43 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1993. ISBN 3-528-07643-7. ISSN 0179-9614. ix + 612 pp. LCCN QA377.I565 1992.

Silvi:2011:QEW

[STFG11] P. Silvi, F. Taddei, R. Fazio, and V. Giovannetti. Quantitative entanglement witnesses of isotropic and Werner classes via local measurements. Journal of Physics A (Mathematical and General), 44(14):145303, 2011. CODEN JPHAC5. ISSN 1751-8113.

Vacanti:2014:TQD

[VFM^{+}14] G. Vacanti, R. Fazio, S. Montangero, G. M. Palma, M. Paternostro, and V. Vedral. Transitionless quantum driving in open quantum systems. New Journal of Physics, 16(May):053017, 2014. CODEN NJOPFM. ISSN 1367-2630.